OpenStudy (anonymous):

How do you tell whether ratios are proportional or not? like 21/18 = 7/6? That's proportional, right? and 23/12 = 7/4, it isn't proportional?

OpenStudy (anonymous):

Proportional means simplified. So 21/18 = 7/6 is proportion. Those two fractions are the same things if put into a calculator. The top and bottom numbers are divided by the same things. 21/3 = 7 and 18/3 = 6